# Polish Academy of Sciences

### v1

@@ 4 Paradigm
Figure 7: Example of swarm of avatars with Chernoff-like faces
@@ 7 Navigating the data
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinhoverde wine database.
Figure 10: Wines data points
@@
In this test Figure 10), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulphates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 Conclusion and Contribution
This paper introduces a new method for visualising high-dimensional data. It uses two natural abilities of the human visual system: understanding of 3D space and recognition of facial features. The method combines a 4D point cloud (rotating in 4D and projected onto a 3D space so as to make its structure accessible to human senses) with expressive icons inspired by Chernoff faces. In this way, it helps people to understand many-dimensional data in a more intuitive and user-friendly way. The 4D cloud can readily be expanded to higher dimensions, but we decided it might overwhelm the user.
The main idea is to divide the data dimensions into two groups. The first group contains features that are easy to understand — such as sensory qualities or user ratings — and these are shown as visual traits of the avatars. The second group includes more abstract or technical parameters, which are shown using 3D positions. If needed, methods such as Principal Component Analysis (PCA) are used to reduce the number of these less intuitive dimensions.
The method has been implemented as a plugin for the open-source platform dpVision. This plugin lets users explore a cloud of avatars in 3D space, where each avatar represents one data point. The user can move around the scene, change the direction of viewing, and adjust the thickness of the visible ”slab” of data, i.e. the extent, in the fourth dimension, of data that will be projected into the 3D viewing space. The assignment of data dimensions to position or appearance can also be changed.
The contribution of this work is threefold. First, it offers a new way to show complex data using both spatial position and visual appearance. Second, it provides an open and working software tool that can be used in practice. Third, it shows the method in action with two datasets: a synthetic cuboid-shaped 4D point cloud and a real-world wine quality dataset with 15 dimensions.
Our experiments show that this approach can help users to better understand complicated data structures. It can be especially useful when not all dimensions are equally important or easy to explain. Since the system is flexible and modular, it can be expanded in the future — for example, by using different avatar types or automatic selection of features.
This method can be helpful in many fields where people need to interpret data visually — such as medicine, quality control, or social sciences — especially when they are not experts in mathematics but still need to make informed decisions.
Our work proposes a new way of thinking about the structure of multidimensional data. Instead of treating all dimensions in the same way, we divide them into two categories: those that are easy to understand and those that are more technical or abstract. This division is then reflected in the visual representation. Dimensions that are more intuitive to the user are shown directly through the appearance of avatars — their shape, size, or facial expression. The other dimensions are either projected into 3D space or reduced using mathematical tools such as PCA.
We believe this distinction is important in many practical situations, where users want to understand what the data means, not just how it is mathematically structured. Unlike traditional visualisation methods, our system provides both spatial and visual channels of information, and assigns them to different types of data in a meaningful way.
To our knowledge, this is the first system that combines avatar-based glyphs with interactive 3D exploration in a way that reflects the human ability to interpret specific features more easily than others. The method has been successfully implemented in the dpVision platform and tested on both synthetic and real-world datasets. This shows that the approach is not only theoretically interesting, but also technically feasible and practically useful.
@@ 9 Future Work
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs — such as neutral geometric shapes or biologically inspired forms like trees — could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user.
@@
A further line of improvement concerns scalability. While the current implementation is effective for datasets with several hundred or a few thousand points, visual clutter becomes an issue when the number of avatars exceeds a certain threshold. Mechanisms such as dynamic filtering, clustering, or zoomable detail views could be added to improve usability for larger datasets.
Finally, a formal evaluation of the system is planned. Although the visualisation method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
Acknowledgements
@@
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Biography Section
Leszek Luchowski (D.Sc.)
has been with the IITiS PAN (formerly ZSAK PAN) since 1983, and part of its Vision Systems Group since 1987. He led the team for a part of this time, and was responsible for a number of research projects in the theory and applications of computer vision. He sustained his PhD thesis in image segmentation in 1994, and his DSc (higher doctorate) in 2014.
Dariusz Pojda (PhD)
specialises in programming (C++, Python, Java and others) and code optimisation. He is interested in data visualisation issues and is the main developer of the dpVision software. He has participated in several research projects on the application of computer vision in fields such as cultural heritage archiving, medicine, and forensics. He has also been involved in work on supporting manufacturing processes using AR/VR systems and applying 3D graphics to visualise quantum computing.

### v2

@@ 5.4 Slab-based Filtering
Figure 8: Part of a sample swarm of avatars with Chernoff-like faces perspectively projected into 3D space.
@@ 7 NAVIGATING THE DATA
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinho Verde wine dataset.
Figure 11: Example of visualization of the data from the Vinho Verde wine dataset.
@@
In this test (Figure 11), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulfates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 CONTRIBUTIONS
This work provides several contributions to the field of multidimensional data visualization:
  1. 1.
    Semantic division of data dimensions
    We propose a novel conceptual framework that divides the dimensions of a dataset into two categories:
    • dimensions that are intuitively understandable to human users (e.g., quality ratings, sensory characteristics),
    • and more abstract or technical dimensions (e.g., chemical or process parameters).
    Unlike previous methods, which typically treat all dimensions equally, we assign intuitive dimensions to visual features of avatars (shape, facial expression, color) and project abstract dimensions into 3D spatial coordinates after possible dimensionality reduction.
  2. 2.
    Integration of avatar-based glyphs with interactive 3D navigation
    While Chernoff faces have historically been used in static 2D representations, we integrate avatar glyphs into a fully interactive 3D environment. Users can freely explore the data cloud, rotate the viewpoint, adjust the visible slab, and examine clusters and patterns dynamically.
  3. 3.
    Practical implementation as an open-source plugin
    We implemented our method as a working plugin for the open-source platform dpVision. The tool supports user-driven mapping of dimensions to visual properties and interactive spatial exploration, offering immediate practical applicability.
  4. 4.
    Demonstration on both synthetic and real-world datasets
    We validated our approach using two datasets:
    • a synthetic 4D anisotropic hypercube (for controlled geometric validation),
    • and the Vinho Verde wine dataset (a 12-dimensional real-world dataset), illustrating the practical benefits of the method in an applied context.
  5. 5.
    Foundations for future extensibility
    The visualization system is designed with modularity and flexibility in mind. Its architecture supports alternative avatar types, adjustable mappings of data dimensions, and dynamic filtering. While this version targets moderately sized datasets, its structure provides a basis for scaling to larger data collections, as further discussed in the Future Work section.
9 CONCLUSION
In this paper, we presented a novel method for visualizing high-dimensional data by combining two natural strengths of the human visual system: the understanding of 3D space and the recognition of facial traits.
Our approach introduces an important distinction between intuitive and abstract data dimensions and maps them differently for optimal interpretability: intuitive features are assigned to avatar characteristics, while abstract features are reduced and projected into 3D space.
We implemented the method as a plugin for the open-source dpVision platform and demonstrated its effectiveness on synthetic and real-world datasets. Preliminary results suggest that this combination of avatar-based glyphs and interactive 3D navigation offers an intuitive and flexible way to explore complex data structures, particularly when users are not domain experts in mathematics.
Further development directions are outlined in the next section and include formal usability studies, exploration of alternative glyph styles to mitigate emotional bias, automated feature mapping, and scalability improvements to handle larger datasets.
By distinguishing between different types of dimensions and tailoring their visual representation, our plugin proposes a human-centered paradigm that could benefit many fields where complex data must be understood and interpreted visually.
@@ 10 FUTURE WORK
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs—such as neutral geometric shapes or biologically inspired forms like trees—could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user. Such alternatives could reduce the emotional bias associated with human-like faces and allow more neutral data interpretation depending on the application domain.
@@
A further line of improvement concerns scalability. While the current implementation is effective for moderately sized datasets (typically up to a few thousand points), its perceptual clarity decreases as the number of avatars grows. Glyph-based methods such as ours are not ideally suited for very large-scale datasets, as visual clutter and cognitive overload quickly become limiting factors. Future versions could integrate mechanisms such as dynamic filtering, clustering, sampling, or zoomable detail views to maintain interpretability in larger data collections. Hybrid solutions that combine glyphs with statistical overviews may also be explored.
Finally, a formal evaluation of the system is planned. Although the visualization method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
11 SUPPLEMENTARY MATERIAL